OpenStudy (anonymous):

Write the slope-intercept equation for the line with the given slope that contains the given point. m=3/2; (7,4)

OpenStudy (dumbcow):

use (y-4) = 3/2(x-7) y-4 = 3/2x - 21/2 y = 3/2x -21/2 +4 y = 3/2x -21/2 + 8/2 y = 3/2x -13/2

OpenStudy (anonymous):

Tks Dc much appreciaTED

OpenStudy (anonymous):

Tks Dc much appreciaTED

OpenStudy (dumbcow):

your welcome :)

OpenStudy (anonymous):

NOW let me ask you this when i solve this problem today i got a different answer would mind looking at it and telling me what i did wrong

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

sry

OpenStudy (dumbcow):

what for?

OpenStudy (anonymous):

for taking so long i closed my eyes for a min and fell asleep i am posting it on the right side now

OpenStudy (dumbcow):

haha ok

OpenStudy (anonymous):

your laughing at me thats not right at all

OpenStudy (anonymous):

y-4=(3/2)(x-7) y-4=(3/2)(4-7) y-4=(1.5)(-3) y-4+4=-1.5+4 y=2.5 2.5=3/2(4)+7